OpenStudy (anonymous):

Mrs. Justin is building a fence for her yard. She has already completed 0.65 of the job. Which fraction below represents the amount of work needed to complete the job?

hartnn (hartnn):

Total work = Work needed to complete + Work already complete 1=x+0.65 x= 1-0.65 =0.35 so is, 35/100 in options ?

OpenStudy (anonymous):

no.....

OpenStudy (anonymous):

what are the multiple choices ?

hartnn (hartnn):

7/20

hartnn (hartnn):

35/100 is also 7/20

OpenStudy (anonymous):

7/20 is corect